OpenStudy (anonymous):

What is the product of 3x^2/x+2 * 5x^2-20/30x^3 ?

mathslover (mathslover):

did u mean this? :

OpenStudy (anonymous):

Yes @mathslover

mathslover (mathslover):

Ok so try to cancel out the terms

mathslover (mathslover):

m i right?

OpenStudy (anonymous):

I think so

mathslover (mathslover):

ok so x^2- 4= (x+2)(x-2) ? right?

OpenStudy (anonymous):

Yessir

mathslover (mathslover):

so this is the answer

OpenStudy (anonymous):

I think it's just x^2-4/2x^2+4x

OpenStudy (anonymous):

Cause that's not one of the answer choices

mathslover (mathslover):

can u name the choices please

OpenStudy (anonymous):

Yes: -2/1+2x X+2/2x X-2/2x

mathslover (mathslover):

ok wait ..

OpenStudy (anonymous):

The one I said before, and those three are the only answer choices

mathslover (mathslover):

well the correct answer is : x-2/2x since

mathslover (mathslover):

we got the answer as : \[\large{\frac{5(x-2)}{10}}\] right?

mathslover (mathslover):